OpenStudy (anonymous):

Given that x<5, rewrite 5x - |x-5| without using absolute value signs

OpenStudy (anonymous):

x-5= x-5 when x>5 5-x when x<5 so 5x-(5-x) =5x-5+x =6x-5

OpenStudy (anonymous):

Ah, so I just "rewrite" the absolute value, then replace the un-needed absolute value signs with parenthesis?

OpenStudy (anonymous):

Yes according to given conditions of x